OpenStudy (anonymous):

is this correct? Determine the equation of the line, in slope-intercept form, that will get your spacecraft from Point B to Point C. (4,8) (6,10) m= y2 - y1 = 10 - 8 = 2 = 1 x2 - x1 = 6 - 4 = 2 y - y1 = m(x - x1) y - 8 = 1(x - 4) y - 8 = x - 4 y = x - 4 + 8 y = x - 12 if not please tell me what i did wrong and how to fix it.

OpenStudy (anonymous):

@Nurali explain please

OpenStudy (nurali):

Sorry u are right.

OpenStudy (anonymous):

@Nurali ok thanks

OpenStudy (nurali):

My Pleasure.

OpenStudy (nurali):

only one mistake y = x - 4 + 8 y=x+4

OpenStudy (isaiah.feynman):

Yes. Also the last line you said "y=x-4+8 gives you y=x-12", sorry that part was wrong (-4+8 = 4) so the equation should y=x+4.

OpenStudy (anonymous):

@Nurali @Isaiah.Feynman thank you

OpenStudy (isaiah.feynman):

No problem ma!

OpenStudy (nurali):

Anytime.